OpenStudy (anonymous):

(-7)^2(-7)= ? and 4^-6 x 4^-1

OpenStudy (zehanz):

The first one looks like (-7)^3=-7*-7*-7 to me. You can do that... Second one:\[4^{-6} \cdot4^{-1}\](a real formula looks better) You have to use these two rules for powers:\[a^b \cdot a^c=a^{b+c}\]and:\[a^{-b}=\frac{ 1 }{ a^b }\]Go ahead, try it!